OpenStudy (anonymous):

an alligator may weigh 350 pounds .a newly hatched alligator weighs only 1over8 , the adult weighs how many times the hatching

OpenStudy (johnweldon1993):

so an alligator weights 350 lbs....a baby weights 1/8 of a pound so \[\frac{ 350 }{ \frac{ 1 }{ 8 } }\] to solve...flip the bottom fraction and turn this into multiplication \[350 \times 8\] this equals??

OpenStudy (anonymous):

okk